OpenStudy (mandii):

explain in complete sentences how adding and subtracting radicals and complex numbers is similar to adding and subtracting polynomials .

OpenStudy (anonymous):

when you add/sub polynomials you combine the like or similar terms based on the variable and exponent. when you add/subtr complex numbers you combine the real parts and the imaginary parts separately. Poly ex: (2+5x) + (3-7x) = 5 -2x Complex ex:(2+5i) + (3-7i) = 5 -2i

OpenStudy (mandii):

Thaank youu !!!